# Life-Code: Central Dogma Modeling with Multi-Omics Sequence Unification

### v1

@@ 3.2 Comparison on Genomic Benchmarks
As DNA sequences are the foundation of the central dogma, we primarily compare with latest DNA foundation models with the SFT evaluation protocol proposed by (Nguyen et al., 2024c; Zhou et al., 2024), following benchmark evaluations for fair comparison. We consider three types of DNA model architectures: (a) classical BERT encoders, including DNABERT (Ji et al., 2021a), DNABERT2 (Zhou et al., 2024), GENA-LM (Fishman et al., 2023), Nucleotide Transformer (NT) (Dalla-Torre et al., 2023), with different tokenizers like VQDNA (Li et al., 2024) and MxDNA (Qiao et al., 2024); (b) RNN-based linear attentions including HyenaDNA (Nguyen et al., 2024c) and Caduceus (Schiff et al., 2024a); (c) CNN model ConvNova (Anonymous, 2025).
@@
We first compare Life-Code with 8 evaluation tasks in the Genomic Benchmark (Grešová et al., 2023), which we adopted results from GenBench implementations (Liu et al., 2024a) and all compared DNA models are fully SFT and reported the top-1 accuracy. Table 2 demonstrates that Life-Code surpasses previous models with remarkable margins since its hybrid designs combine the advantages of both RNN and self-attention blocks.
@@
We then evaluate Life-Code with 24 long-range genomic tasks on the GUE benchmark (Zhou et al., 2024), as shown in Table 3, where 7 widely adopted practical genomic tasks are conducted, including Epigenetic Mark Prediction (EMP) for Yeast, Transcription Factor Prediction on Mouse and Human genome, Promoter Detection, Core Promoter Detection, Splice Site Prediction, and Covid Variants Classification. All results are adopted from DNABERT2 or each paper with their official SFT setups, and Matthews Correlation Coefficient (MCC) is reported as the metric. Compared to popular DNA models, our Life-Code shows superiority in gene annotation tasks, which are important regulation processes in the central dogma, while achieving competitive results in Epigenetic regulating tasks.
Table 4: Comparison on mRNA Splicing Site Prediction with two splicing datasets from SpliceAI and Spliceator. Top-1 AUC score or F1 score are reported for DNA models or RNA models with SFT evaluation, respectively. The best and the second best results are marked as the bold and underlined types.
@@
Spliceator 91.0 91.9 92.1 94.8 94.5
3.3 Comparison on RNA Splicing Site Prediction
Since precursor messenger RNAs (pre-mRNAs) splicing is a crucial process in eukaryotic gene expression, we then evaluate Life-Code on Splicing Site Prediction tasks as binary classification tasks. We consider two typical datasets with the SFT evaluation protocol. As for SpliceAI dataset (Jaganathan et al., 2019), we evaluate DNA models with top-1 AUC scores and find that Life-Code consistently outperforms popular SpliceAI and DNA models, indicating its great ability to model the central dogma. As for Spliceator dataset (Scalzitti et al., 2021), we then consider RNA splicing methods (Spliceator, DDSP (Naito, 2019), RINALMo (Penic et al., 2024)) and RNA language model, RNA-FM (Shen et al., 2024), with RNA data of multi-species. As shown in Table 4, our Life-Code achieves competitive F1 scores with the specially designed splicing method and surpasses RNA-FM.
Table 5: Comparison on Zero-shot Protein Fitness Prediction. SRCC (%) is reported on DMS datasets. The best and the second best results are marked as the bold and underlined types. ESM2 is marked as gray since it takes the amino acids as the input, while other models utilize the gene sequences.
Method GenSLM NT-2500M EVO ESM2-650M Life-Code
# Params 2.5B 2.5B 7B 650M 350M
Bacteria 24.7 9.4 45.3 51.2 45.7
Human 6.9 4.7 11.1 37.5 22.4
@@ 3.4 Comparison on Protein Fitness Prediction
Following EVO (Nguyen et al., 2024a), we then evaluate foundation models pre-trained with DNA or protein data with Protein Fitness Prediction tasks using zero-shot evaluation protocol (without task-specific SFT). Deep Mutational Scanning (DMS) datasets (Notin et al., 2022; Nguyen et al., 2024a) introduce a comprehensive set of mutations into protein CDS to measure their effects on fitness, where fitness serves as a metric for how effectively a protein performs a specific function. As shown in Table 5, bacterial (prokaryote) and human (eukaryote) protein fitness is evaluated by Spearman’s Rank Correlation Coefficient (SRCC), where the protein specialist model ESM2 achieves the best result. Life-Code outperforms other DNA models like GenSLM (Zvyagin et al., 2022) and EVO-7B with fewer parameters, which might be attributed to its modeling of the translation and folding processes in the central dogma.
Table 6: Comparison on Multi-omics Tasks. Top-1 accuracy (%) is reported for ncRNA-Protein Interaction with SFT and few-shot central dogma evaluation. Bold type denotes the best result.
@@
Life-Code 93.45 85.63
3.5 Comparison on Multi-omics Tasks
We conduct further comparisons with proposed biological foundation models (EVO (Nguyen et al., 2024a) and LucaOne (He et al., 2024)) with multi-omics tasks. Following LucaOne, we first evaluate the non-coding RNA and protein interactions (ncRPI) task using SFT evaluation, which aims to predict the interaction strengths between ncRNAs and proteins. As shown in Table 6, Life-Code could take cDNA and Amino Acid sequences in the unified input format and achieve the competitive result as the supervised LucaOne, which should process two inputs with separate modules. Then, we evaluate the modeling of transcription and translation processes with binary classification tasks between cDNA and the protein sequences, where Life-Code achieves the best accuracy over existing models.
@@ 3.5 Comparison on Multi-omics Tasks
Life-Code 350M 62.2 87.6 99.3 99.4 99.5
@@ 3.6 Empirical Analysis
After verifying the effectiveness with extensive comparison results, we further consider long-sequence tasks like HyenaDNA (Nguyen et al., 2024c) in addition to Figure 7. We compare the inference throughput of our bi-directional hybrid model to baselines (FlashAttention and HyenaDNA). In Table 7, empirical profiling shows that LifeCode achieves competitive speed with these specialized methods while retaining strong global modeling capacity. In particular, mixing a small fraction of MHSA enables better downstream performance without incurring a large computational penalty, e.g., splicing site prediction in Table 4.
@@ 4 Related Works
Recent years have seen rapid progress in DNA foundation models that leverage Transformer-like architectures. Early attempts such as DNABERT (Ji et al., 2021b) treated genomic sequences akin to linguistic tokens, enabling pre-trained contextual representations. Follow-up work DNABERT-2 (Zhou et al., 2023) expanded this paradigm by introducing more efficient training protocols and supporting multi-species data. In parallel, The Nucleotide Transformer (Dalla-Torre et al., 2023) demonstrated the feasibility of scaling up Transformer architectures for human genomics. Beyond Transformers, kernel-based or hierarchical approaches emerged—e.g., HyenaDNA (Nguyen et al., 2024b), which reduces the quadratic complexity of attention for extremely long sequences. Meanwhile, alternative designs like Caduceus (Schiff et al., 2024b) incorporate selective structural priors for long-range DNA modeling, and MxDNA (Qiao et al., 2024) explores adaptive tokenization schemes that automatically discover suitable patterns for genomic data.
@@
On the protein side, AlphaFold series (Jumper et al., 2021a, b) revolutionized structure prediction, catalyzing a surge in protein-based language models such as ESM-2 (Lin et al., 2022), which refines large-scale protein embeddings. Beyond proteins, RNA modeling also gained traction; for instance, (Shen et al., 2024) employed a language-model-based technique for accurate 3D structure prediction. These efforts underscore the trend toward specialized architectures for each biomolecule yet also highlight the desire for integrative multi-omics solutions.
@@
Recent studies aim to bridge the gap between genomic and protein sequences, pushing beyond single-modality tasks. CD-GPT (Zhu et al., 2024) explicitly connects DNA, RNA, and proteins through the central dogma, while BSM (Xiang et al., 2024) highlights the potential for small but effective models covering genes and proteins simultaneously. Evo (Nguyen et al., 2024a) similarly integrates molecular and genome-scale data with a generalized sequence-modeling approach. Closely related is LucaOne (He et al., 2024), which advocates a unified nucleic acid and protein language for biological representation.
@@ 5 Conclusion and Limitations
We have introduced Life-Code, a method that unifies DNA, RNA, and protein sequences by mapping the latter two back to a DNA-like representation. This design offers straightforward multi-omics integration and maintains biological interpretability via differential encoding of coding versus non-coding regions. To handle the long-range dependencies inherent in genomic data, we employ an efficient symmetric convolution-based architecture. Moreover, knowledge distillation from large protein models allows our approach to scale without excessive resource demands.
@@
Impact Statements
This paper presents a novel approach, Life-Code, aimed at advancing the field of machine learning and bioinformatics by unifying the modeling of multi-omics data through the guidance of the Central Dogma of molecular biology. By integrating DNA, RNA, and protein data into a single framework, our work has the potential to improve the understanding of complex biological processes and drive advancements in genomics, proteomics, and systems biology.
Ethical Aspects
This work has significant potential applications in fields such as personalized medicine, drug discovery, and disease diagnostics. However, we acknowledge the need to address concerns related to data privacy and the responsible use of sensitive genetic information. We have adhered to ethical guidelines in our experiments, ensuring that no private or identifiable data was used.
@@
Figure A1: Data Collection Pipeline. We collect the reference sequences from the NCBI RefSeq database as the DNA dataset and collect the cDNA of coding sequences with the corresponding amino acids from the GenBank and UniRef50 databases to build up our DNA-AA paring datasets.
Appendix A Implementation Details
@@ A.1 Pre-training Dataset
We collect two datasets for pre-training of Life-Code models, i.e., a pure DNA dataset and a DNA-AA pairing dataset, with the collection process shown in Figure A1. As for the DNA dataset, we collect the reference sequences (RefSeq) of multiple species to ensure generalization abilities from the database of the National Center for Biotechnology Information (NCBI) at https://www.ncbi.nlm.nih.gov, following the Multi-species Genomes [^1] Multi-species Genomes are originally provided in https://huggingface.co/datasets/InstaDeepAI/multi_species_genomes, which is further extended by DNABERT-2 in https://github.com/MAGICS-LAB/DNABERT_2 provided by Nucleotide Transformer (Dalla-Torre et al., 2023) and DNABERT2 (Zhou et al., 2024). As for the DNA-AA pairing dataset, we collect the cDNA of coding sequences (CDS) and its corresponding Amino Acids (AA) in the GenBank database at https://www.ncbi.nlm.nih.gov/genbank, which aims to model the transcription and translation processes of the central dogma. We also collect some Amino Acids in the UniRef50 database following LucaOne (He et al., 2024) and obtain their corresponding cDNA by reverse translation with online tools. We provide detailed information for used datasets in Table A1.
@@ A.2 Life-Code Tokenizer
There are two vocabularies used in Life-Code. The unified vocabulary only uses 4 nucleotides {A, T/U, C, G} of nucleic acid with 5 special tokens, including “[U]”/“[UNK]”, “[PAD]”, “[CLS]”, “[SEP]”, and “[MASK]” for unknown nucleotides, padding tokens, the class token, separator tokens, and masking tokens. Meanwhile, the Life-Code can also use the codon vocabulary (i.e., the 3-mer of 4 nucleotides that constructs 64 codon tokens), which could be merged into 20 amino acids of protein (20 uppercase letters excluding “B”, “J”, “O”, “U”, “X”, and “Z”). It can only be applied when the length of an input sequence is multiples of 3, i.e., the cDNA of amino acids or matured mRNA (CDS). The pre-trained protein language model employs the amino acid vocabulary of ESM-2 (Lin et al., 2022).
@@
Base learning rate $2\times 10^{-4}$ $1\times 10^{-4}$
Batch size 512 256
@@
As shown in Figure 4, we pre-train the Life-Code tokenizer with the DNA de-tokenizer and Amino Acid de-tokenizer by AdamW optimizer for 100,000 iterations (randomly sampled datasets) with a basic learning rate of 2e-4 and a batch size of 512, as detailed in Table A2. We utilize 8 Nvidia A100-80G GPUs with a per-GPU batch size of 8 and a gradient accumulation time of 4.
@@ A.3 Life-Code Encoder
As shown in Table A2, the Life-Code encoder has 24 layers in total with the embedding dim of 1024 with the following designs: (1) Mixture of GDN blocks (DeltaNet (Yang et al., 2025)) and multi-head self-attention (MHSA) blocks as a hybrid model, especially every 11 GDN blocks followed by a self-attention block like MiniMax-01 (MiniMax et al., 2025), which could utilize the complementary properties of GDN and MHSA while maintaining efficiency. (2) The model macro design employs pre-norm (Wang et al., 2019a) with RMSNorm (Zhang & Sennrich, 2019), Layer Scale (Touvron et al., 2021), Rotary Position Embedding (RoPE) (Su et al., 2021), SwiGLU (Touvron et al., 2023), and FlashAttention implementations to facilitate training large-scale models stably with long sequences. (3) During pre-training, we apply the packing strategy (Warner et al., 2024) to build up a long sequence with several CDS, which compromises the gap between different lengths of the reference sequence and the coding sequences, as shown in Figure 3.
@@
As shown in Figure 5, we further pre-train the Life-Code tokenizer and Encoder with three tasks in Eq. 5 for 1M steps with the batch size of 256 and the basic learning rate of $1\times 10^{-4}$. We adopt 15% random masking in BERT for Masked DNA Reconstruction and the 3-mer span masking for CDS-to-Amino-Acid Translation. As for Knowledge Distillation from a pre-trained Protein LM, we adopted ESM2-650M (esm2_t33_650M_UR50D(Lin et al., 2022) and a protein decoder with the output dimension of 1280. During the warmup periods, the maximum sequence length is 1024, with a linear warmup of the learning rate for 10 iterations. After that, the maximum sequence length is set to 4k with the learning rate adjusted by the Cosine Annealing scheduler (decay to $1\times 10^{-6}$). We utilize 8 Nvidia A100-80G GPUs with a per-GPU batch size of 2 and a gradient accumulation time of 16.

### v2

@@ 3.2 Comparison on Genomic Benchmarks
As DNA sequences are the foundation of the central dogma, we primarily compare with the latest DNA foundation models with the SFT evaluation protocol proposed by (nips2024hyenadna, ; iclr2024dnabert2, ), following benchmark evaluations for fair comparison. We consider three types of DNA model architectures: (a) classical BERT encoders, including DNABERT (BioInfo2021dnabert, ), DNABERT2 (iclr2024dnabert2, ), GENA-LM (fishman2023GenaLM, ), Nucleotide Transformer (NT) (NM2023NucleotideTrans, ), with different tokenizers like VQDNA (icml2024vqdna, ) and MxDNA (nips2024MXDNA, ); (b) RNN-based linear attentions including HyenaDNA (nips2024hyenadna, ) and Caduceus (icml2024caduceus, ); (c) CNN model ConvNova (iclr2025convnova, ).
@@
We first compare Life-Code with 8 evaluation tasks in the Genomic Benchmark (BMC2023genomicbenchmark, ), which we adopted results from GenBench implementations (liu2024genbench, ), and all compared DNA models are fully SFT and reported the top-1 accuracy. Table 1 demonstrates that Life-Code surpasses previous models with remarkable margins since its hybrid designs combine the advantages.
@@
We then evaluate Life-Code with 24 long-range genomic tasks on the GUE benchmark (iclr2024dnabert2, ), as shown in Table 2, where 7 widely adopted practical genomic tasks are conducted, including Epigenetic Mark Prediction (EMP) for Yeast, Transcription Factor Prediction on Mouse and Human genome, Promoter Detection, Core Promoter Detection, Splice Site Prediction, and Covid Variants Classification. All results are adopted from DNABERT2 or each paper with their official SFT setups, and the Matthews Correlation Coefficient (MCC) is reported as the metric.
3.3 Comparison on RNA Splicing Site Prediction
Since precursor messenger RNAs (pre-mRNAs) splicing is a crucial process in eukaryotic gene expression, we then evaluate Life-Code on Splicing Site Prediction tasks as binary classification tasks. We consider two typical datasets with the SFT evaluation protocol. As for the SpliceAI dataset (JAGANATHAN2019SpliceAI, ), we evaluate DNA models with top-1 AUC scores and find that Life-Code consistently outperforms popular SpliceAI and DNA models, indicating its great ability to model the central dogma. As for Spliceator dataset (BMC2021spliceator, ), we then consider RNA splicing methods (Spliceator, DDSP (naito2019DDSP, ), RINALMo (Penic2024RiNALMo, )) and RNA language model, RNA-FM (shen2024rnafm, ), with RNA data of multi-species. As shown in Table 4, our Life-Code achieves the best F1 scores with the specified baselines.
Table 3: Comparison on Zero-shot Protein Fitness Prediction. SRCC (%) is reported on DMS datasets.
Method GenSLM NT-2500M EVO Life-Code
# Params 2.5B 2.5B 7B 500M
Bacteria 24.7 9.4 45.3 46.4
Human 6.9 4.7 11.1 23.8
Table 4: Comparison on mRNA Splicing Site Prediction with two splicing datasets from SpliceAI and Spliceator. The top-1 AUC score is reported.
@@ 3.3 Comparison on RNA Splicing Site Prediction
Spliceator 91.0 91.9 92.1 94.8 95.0
@@ 3.4 Comparison on Protein Fitness Prediction
Following EVO (nguyen2024evo, ), we evaluate foundation models in a strict zero-shot setting on Protein Fitness Prediction. Deep Mutational Scanning (DMS) datasets (icml2022Tranception, ; nguyen2024evo, ) introduce exhaustive mutations into protein coding sequences (CDS) and report fitness scores that reflect functional efficacy. For both bacterial and human proteins, we compute Spearman’s rank correlation (SRCC) between model scores and experimental measurements (Table 3). The protein-specialist ESM-2 sets the upper bound, yet Life-Code surpasses all DNA-based baselines including GenSLM (ijhpca2023genslms, ) and the 7-B-parameter EVO model—despite using fewer parameters. We attribute this advantage to Life-Code’s unified pre-training on DNA, RNA, and amino acids, which captures translation constraints and folding cues along the central-dogma pathway, enabling strong generalization from nucleotide space to protein fitness without any task-specific fine-tuning.
Table 5: Comparison on Multi-omics Tasks. Top-1 accuracy (%) is reported for ncRNA-Protein Interaction with SFT and few-shot central dogma evaluation.
@@
Life-Code 94.12 86.51
@@
Life-Code 500M 62.8 88.1 99.3 99.4 99.5
Figure 8: Model accuracy vs. throughput in K tokens/s for different hybrid layer ratios. A shift from 11:1 to 5:1 slightly slows throughput with minimal accuracy loss; 3:1 exceeds memory limits.
3.5 Comparison on Multi-omics Tasks
We conduct further comparisons with proposed biological foundation models (EVO (nguyen2024evo, ) and LucaOne (he2024lucaone, )) with multi-omics tasks. Following LucaOne, we first evaluate the non-coding RNA and protein interactions (ncRPI) task using SFT evaluation, which aims to predict the interaction strengths between ncRNAs and proteins. As shown in Table 5, Life-Code could take cDNA and Amino Acid sequences in the unified input format and achieve a competitive result as the supervised LucaOne, which should process two inputs with separate modules. Then, we evaluate the modeling of transcription and translation processes with binary classification tasks between cDNA and protein sequences, where Life-Code achieves the best accuracy over existing models.
@@ 3.6 Empirical Analysis
Figure 8 shows that a pure RNN stack (12:0) is fastest ( $\sim$45 K token/s) but sacrifices 2–3 pp of accuracy. Injecting just one MHSA layer (11:1) restores nearly all lost performance—DNA 95 to 97 %, Protein 32 to 34 %, CD 83 to 86 %—while reducing throughput by only $\sim$9 %. A 5:1 mix boosts accuracy further with a modest extra cost, whereas 3:1 exhausts memory. Thus, adding a small fraction of MHSA markedly improves downstream accuracy without a large computational penalty.
@@ 4 Related Works
Recent DNA foundation models have evolved swiftly from DNABERT ji2021dnabert , which first treated genomic 6-mers as linguistic tokens, to DNABERT-2 zhou2023dnabert2 and the large-scale Nucleotide Transformer NM2023NucleotideTrans . To tame quadratic attention on longer sequences, HyenaDNA nguyen2024hyenadna replaces full self-attention with efficient hierarchical kernels. Other variants embed biological priors—Caduceus schiff2024caduceus for long-range structure and MxDNA nips2024MXDNA for adaptive tokenization—illustrating a broad shift toward specialized, scalable architectures for genomics.
@@
AlphaFold jumper2021alphafold ; jumper2021alphafold2 ignited the protein-LM wave, inspiring models like ESM-2 lin2022ESM2 . RNA efforts followed, e.g., RNA-FM shen2024rnafm for 3D structure prediction. While these biomolecule-specific architectures excel individually, recent work now fuses DNA, codon, and protein signals prakash2024bridging ; bioinfo2024genomic ; BioInfo2023NTA ; nmil2024CaLM , delivering notable downstream gains Ren2024COMET and pointing toward unified multi-omics models.
@@
Recent studies aim to bridge the gap between genomic and protein sequences, pushing beyond single-modality tasks. CD-GPT zhu2024CDGPT explicitly connects DNA, RNA, and proteins through the central dogma, while BSM xiang2024BSM highlights the potential for small but effective models covering genes and proteins simultaneously. EVO nguyen2024evo similarly integrates molecular and genome-scale data with a generalized sequence-modeling approach. Closely related is LucaOne he2024lucaone , which advocates a unified nucleic acid and protein language for biological representation.
@@ 5 Conclusion and Limitations
Life-Code unifies DNA, RNA, and protein by remapping all sequences to nucleotides, enabling seamless multi-omics integration while preserving biological meaning (coding vs non-coding). A symmetric convolution backbone captures long-range genomic context, and distilling knowledge from large protein LMs keeps training efficient.
@@
Figure A1: Data Collection Pipeline. We collect the reference sequences from the NCBI RefSeq database as the DNA dataset and collect the cDNA of coding sequences with the corresponding amino acids from the GenBank and UniRef50 databases to build up our DNA-AA pairing datasets.
@@ A.1 Pre-training Dataset
We collect two datasets for pre-training of Life-Code models, i.e., a pure DNA dataset and a DNA-AA pairing dataset, with the collection process shown in Figure A1. As for the DNA dataset, we collect the reference sequences (RefSeq) of multiple species to ensure generalization abilities from the database of the National Center for Biotechnology Information (NCBI) at https://www.ncbi.nlm.nih.gov, following the Multi-species Genomes [^1] Multi-species Genomes are originally provided in https://huggingface.co/datasets/InstaDeepAI/multi_species_genomes, which is further extended by DNABERT-2 in https://github.com/MAGICS-LAB/DNABERT_2 provided by Nucleotide Transformer [8] and DNABERT2 [64]. As for the DNA-AA pairing dataset, we collect the cDNA of coding sequences (CDS) and its corresponding Amino Acids (AA) in the GenBank database at https://www.ncbi.nlm.nih.gov/genbank, which aims to model the transcription and translation processes of the central dogma. We also collect some Amino Acids in the UniRef50 database following LucaOne [13] and obtain their corresponding cDNA by reverse translation with online tools. We provide detailed information for used datasets in Table A1.
@@ A.2 Life-Code Tokenizer
There are two vocabularies used in Life-Code. The unified vocabulary only uses 4 nucleotides {A, T/U, C, G} of nucleic acid with 5 special tokens, including “[U]”/“[UNK]”, “[PAD]”, “[CLS]”, “[SEP]”, and “[MASK]” for unknown nucleotides, padding tokens, the class token, separator tokens, and masking tokens. Meanwhile, the Life-Code can also use the codon vocabulary (i.e., the 3-mer of 4 nucleotides that constructs 64 codon tokens), which could be merged into 20 amino acids of protein (20 uppercase letters excluding “B”, “J”, “O”, “U”, “X”, and “Z”). It can only be applied when the length of an input sequence is multiples of 3, i.e., the cDNA of amino acids or matured mRNA (CDS). The pre-trained protein language model employs the amino acid vocabulary of ESM-2 [26].
@@
Base learning rate $4\times 10^{-4}$ $1\times 10^{-4}$
Batch size 1024 256
@@
As shown in Figure 4, we pre-train the Life-Code tokenizer with the DNA de-tokenizer and Amino Acid de-tokenizer by AdamW optimizer for 100,000 iterations (randomly sampled datasets) with a basic learning rate of $2\times 10^{-4}$ and a batch size of 1024, as detailed in Table A2. We utilize 8 Nvidia A100-80G GPUs with a per-GPU batch size of 8 and a gradient accumulation time of 16.
@@ A.3 Life-Code Encoder
As shown in Table A2, the Life-Code encoder has 24 layers in total with the embedding dim of 1024 with the following designs: (1) Mixture of GDN blocks (DeltaNet [60]) and multi-head self-attention (MHSA) blocks as a hybrid model, especially every 11 GDN blocks followed by a self-attention block like MiniMax-01 [32], which could utilize the complementary properties of GDN and MHSA while maintaining efficiency. Notice that the DeltaNet variants [59] introduce short-term depth-wise convolution after the query/key/value projections in the token mixing block, which requires slightly more parameters than the standard Transformer block. (2) The model macro design employs pre-norm [55] with RMSNorm [61], Layer Scale [52], Rotary Position Embedding (RoPE) [51], SwiGLU [53], and FlashAttention implementations to facilitate training large-scale models stably with long sequences. (3) During pre-training, we apply the packing strategy [57] to build up a long sequence with several CDS, which compromises the gap between different lengths of the reference sequence and the coding sequences, as shown in Figure 3.
@@
As shown in Figure 5, we further pre-train the Life-Code tokenizer and Encoder with three tasks in Eq. 5 for 1M steps with a batch size of 256 and a basic learning rate of $1\times 10^{-4}$. We adopt 15% random masking in BERT for Masked DNA Reconstruction and the 3-mer span masking for CDS-to-Amino-Acid Translation. As for Knowledge Distillation from a pre-trained Protein LM, we adopted ESM2-650M (esm2_t33_650M_UR50D[26] and a protein decoder with the output dimension of 1280. During the warmup periods, the maximum sequence length is 1024, with a linear warmup of the learning rate for 10 iterations. After that, the maximum sequence length is set to 4k with the learning rate adjusted by the Cosine Annealing scheduler (decay to $1\times 10^{-6}$). We utilize 8 Nvidia A100-80G GPUs with a per-GPU batch size of 2 and a gradient accumulation time of 16.